PE of molecular weight = 4.9×105gm/mol has repeat unit = −(CH2)−. Calculate the number of repeat units
1
Expert's answer
2020-04-06T09:19:06-0400
"N=\frac{M}{M_{CH_2}}"

"N=\frac{4.9\cdot 10^5}{14}=35000"